OpenStudy (anonymous):

To factor 9x^2-16 you can first rewrite the expression as: A.(3x^2)-(4)^2 B.(3x-4)^2 C.(3x)^2-(4) D. None of the above

OpenStudy (usukidoll):

we have perfect square numbers 9 and 16 a perfect square is the product of a rational number multiplied by itself. the product of a polynomial multiplied by itself. We have the formula \[(a^2-b^2) = (a+b)(a-b) \]

OpenStudy (usukidoll):

A perfect square is a number that can be expressed as the product of two equal integers.

OpenStudy (anonymous):

(3x+4)(3x−4)

OpenStudy (welshfella):

this is called the difference of 2 squares and is described by the identity UsukiDoll posted a^2 - b^2 = (a + b)(a - b)

OpenStudy (usukidoll):

I forgot about typing "difference of 2 squares". thank you for pointing that out @welshfella